OpenStudy (anonymous):

when I factor y^3-4y the answer is y(y^2+2)(y-2) right

ganeshie8 (ganeshie8):

can you show how did u arrive at that

OpenStudy (anonymous):

y(y^2+2)(y-2) (y^2-2y+2y_4) y(y^2-4) =y^3-4y

OpenStudy (anonymous):

y(y^2+2)(y-2) y(y^2-2y+2y-4) y(y^2-4) =y^3-4y

ganeshie8 (ganeshie8):

ok can you tell me this : \(y^2 \times y = ?\)

OpenStudy (anonymous):

y^3

ganeshie8 (ganeshie8):

good :) , then how did you get the second step in your solution ?

OpenStudy (anonymous):

worked whats in the () first the I combined like terms then i distributed the y that was on the out side of the () so did i work the problem right? is that the right factor

ganeshie8 (ganeshie8):

its not correct

ganeshie8 (ganeshie8):

you did a mistake at second step

ganeshie8 (ganeshie8):

you wrote y^2 x y as y^2, instead of y^3

ganeshie8 (ganeshie8):

check it once